# Multispectral airborne laser scanning for tree species classification: a benchmark of machine learning and deep learning algorithms

### v1

@@ 4 Results and Discussion
In this section, we evaluate the performance of the studied species classification methods on the multispectral data collected with the HeliALS and Optech Titan systems. First in Sec. 4.1, we present the results of the benchmarking contest and discuss key factors affecting the classification accuracy, such as spectral information, the use of a deep learning or a machine learning model, model architecture and training strategy. In Sec. 4.2, we analyze the scaling behavior of classification accuracy for selected deep learning and machine learning models as a function of training set size and point density to study the impact of these factors in more detail and to highlight the differences of deep learning and machine learning methods. Subsequently, we investigate the effect of segmentation quality and crown class on species classification in Sec. 4.3. Finally, we compare our results to previous studies and provide further discussion of our results in Sec. 4.4.
@@ 4.1 International benchmarking of species classification methods
Figure 6: (a) Comparison of overall (filled bar) and macro-average (outline) accuracy for methods using HeliALS data, with point-based DL methods shown with dark blue color, image-based DL methods shown with light blue color, and machine learning methods shown with red color. SA-Detailview-DL-2D was not retrained on the training set of the current study, and the model parameters were directly transferred from Puliti et al. (2025). When neglecting aspens, rowans and alders not present in the FOR-species20K dataset (Puliti et al., 2025), SA-Detailview-DL-2D reached an overall accuracy of 58.8% and a macro-average accuracy of 45.4%. (b) Same as (a) but for the methods using the Optech Titan dataset.
@@
The species classification methods introduced in Sec. 3.2 were benchmarked using overall accuracy, macro-average accuracy, species-wise precision and recall as the key metrics. We compare the overall and macro-average accuracy for the methods using the HeliALS data in Fig. 6(a) and for the methods using the Optech Titan data in Fig. 6(b). In general, the methods using the HeliALS data reached a higher overall and macro-average accuracy compared to methods using the Optech Titan data, which can be attributed to an order of magnitude higher point density and improved point cloud quality provided by the HeliALS system, see Sec. 2.2.
@@
For the HeliALS dataset, FGI-PointTransformer-DL-3D reached the highest overall accuracy of 87.9%, and the corresponding method with a weighted loss function, FGI-PointTransformerWeighted-DL-3D, achieved the highest macro-average accuracy of 74.5%. In addition to the point transformer model, the other best performing methods on the HeliALS dataset were also point-based deep learning methods, with FGI-DGCNN-DL-3D and FGI-PointNet-DL-3D reaching an overall (macro-average) accuracy of 87.0% (67.3%) and 85.9% (64.2%), respectively. The best image-based deep learning method, DetailView-DL-2D, reached an overall (macro-average) accuracy of 84.3% (63.9%) and the only machine learning model, FGI-RF-ML, had an overall (macro-average) accuracy of 83.2% (61.3%). For most of the methods, the macro-average accuracy was significantly lower than the overall accuracy due to the inherent difficulty of predicting the minority species. The point transformer model clearly outperformed the other methods in the prediction of the minority species, which resulted in a significant improvement of the macro-average accuracy, especially when using the weighted loss function.
@@
Figure 7: (a) Normalized confusion matrix of FGI-RF-ML obtaining the highest macro $F_{1}$ score (0.64) among the maching learning models using HeliALS data. FGI-RF-ML produced 55 missing predictions. (b) Normalized confusion matrix of FGI-PointTransformerWeighted-DL-3D obtaining the highest macro $F_{1}$ score (0.73) among the deep learning models using HeliALS data. (c) Normalized confusion matrix of IBL-BalancedRF-ML obtaining the highest macro $F_{1}$ score (0.57) among the machine learning models using Optech Titan data. (b) Normalized confusion matrix of FGI-PointTransformer-DL-3D obtaining the highest macro $F_{1}$ score (0.58) among the deep learning models using Optech Titan data. In all panels, the matrix elements have been normalized with the row sums. Thus, diagonal values represent species-wise recalls. Precisions are displayed below the matrix. Row and column sums are shown to the right and bottom of the matrix in gray.
@@
For the Optech Titan dataset, the best performing method was a random forest classifier, IBL-BalancedRF-ML, which reached an overall accuracy of 79.9% and a macro-average accuracy of 57.6%. The point transformer method, FGI-PointTransformer-DL-3D, ranked second with an overall accuracy of 79.6% and a macro-average accuracy of 56.0%. For a larger training set size, the deep learning method is expected to outperform the machine learning model as discussed in Sec. 4.2. On the Optech Titan dataset, most of the participating methods were RF classifiers or their variants. The best and worst-performing RF models had a difference of 7.9 (10.3) percentage points in terms of the overall (macro-average) accuracy implying that the performance of the various RF approaches was relatively consistent. This suggests that RF is a robust and reliable method, providing reasonable accuracy across different implementations. Based on the submissions by UEF, a RF classifier also outperformed both LGBM and SVM classifiers, though the difference was only a few percentage points.
@@
It is difficult to draw exhaustive conclusions on the factors causing the performance differences between the studied RF classifiers. However, the method descriptions in Sec. 3.2.3 suggest that the best-performing RF classifiers often utilized multi-channel intensity and echo features as well as various histogram features, such as intensity histograms. Importantly, intensity features appear to be more important than geometric features since the second best RF method, UEF-RF-ML, only utilized a single geometric feature, but still managed to reach a good performance thanks to various intensity and echo features. When it comes to training strategies, the best performing model, IBL-BalancedRF-ML, cleaned the training dataset by removing under- or over-segmented trees, and applied a balanced RF procedure (Chen et al., 2004) to up-sample the number of training segments for minority species. However, LUKE-MultiRF-ML and Aalto-RF-ML also applied strategies to mitigate class imbalance but they were outperformed by several methods that did not use any strategy for class imbalance mitigation. Some RF methods had trouble evaluating all of their features on the sparsest segments, thus hampering their prediction. On sparse ALS data, it is therefore advisable to use sufficiently simple geometric features. Otherwise, the probability of missing feature values may increase, and imputation strategies are needed.
@@
Figure 8: Precision and recall values for the best machine learning and deep learning methods for HeliALS (Panels (a) and (b)) and Optech Titan (Panels (c) and (d)) datasets. Best methods were chosen based on the macro-average $F_{1}$ score. Median values for precision and recall were calculated across the methods utilizing the same dataset and approach (DL/ML). (a) Precision and recall values for FGI-RF-ML utilizing HeliALS data (Macro $F_{1}=0.64$) (colored bars). (b) Precision and recall values for the best DL method, FGI-PointTransformerWeighted-DL-3D, (Macro $F_{1}=0.73$) (colored bars) and the median across other DL methods (dashed line bars) on the HeliALS dataset. (c) Precision and recall values for the best ML method (Macro $F_{1}=0.57$) (colored bars) and the median across other ML methods (dashed line bars) on the Optech Titan data. (d) Precision and recall values for the best DL method (Macro $F_{1}=0.58$) (colored bars) and the median across other DL methods on the Optech Titan data (dashed line bars).
@@
Using the HeliALS data, we also tested a sensor-agnostic method, SA-Detailview-DL-2D. The method had been previously trained on the FOR-species20K dataset  (Puliti et al., 2025), and we did not re-train it on the current dataset as explained in Sec. 3.2.2. The method did not utilize any spectral information either. On the full test set, SA-Detailview-DL-2D achieved an overall (macro-average) accuracy of 51.4% (30.3%) whereas the overall (macro-average) accuracy was 58.8% (45.4%) when neglecting aspen, rowan, and alder not present in the FOR-species20K dataset. Importantly, the performance of the DetailView model was greatly improved by training the model from scratch using the training dataset of our study. This resulted in an overall (macro-average) accuracy of 84.3% (63.9%) for DetailView-DL-2D. The finding suggests that the model architecture itself is promising for species classification, but the model trained on the global FOR-species20K dataset did not generalize well enough for our ALS dataset. Possible reasons behind the poor out-of-domain performance of SA-Detailview-DL-2D are the relatively small proportion of high-density ALS data compared to TLS and MLS data in the FOR-species20K training dataset, the impact of segmentation artifacts, and the poor representativeness of the species of this study in the FOR-species20K training dataset.
@@
To understand differences in the classification accuracy between different species, we present confusion matrices of the best-performing deep learning and machine learning models for both the HeliALS and Optech Titan datasets in Figs. 7(a)–(d). Furthermore, Figures 8(a)–(d) illustrate the species-wise precision and recall for the best methods on both datasets together with the median of the machine learning or deep learning approaches. Based on Figs. 7(a)–(d), we observe that pines, spruces, birches, and maples are reliably classified on both datasets across the models. The other deciduous trees representing minority species (see Fig. 3) are associated with significantly lower species-wise recall and precision. Based on the confusion matrices, we observe that aspens are often classified as birches, rowans as maples, lindens as maples, alders as aspens, and oaks as all other deciduous trees across the different models and datasets. The difficulty in predicting these species arises partly due to their low frequency in the benchmarking dataset and partly due to a rather similar visual appearance as further discussed in Sec. 4.4. For the Optech Titan dataset, we observe no clear qualitative difference between the predictions of the best deep learning and the best machine learning model. The dense HeliALS dataset enables a more accurate classification of the minority species. Importantly, the best deep learning method, FGI-PointTransformerWeighted-DL-3D, clearly outperforms the best machine learning method FGI-RF-ML in the prediction of minority species on the HeliALS dataset, which highlights the promise of deep learning for species classification on dense ALS point clouds. Importantly, FGI-PointTransformerWeighted-DL-3D achieved a recall of 76.9% and a precision of 66.3% for aspen that has an important role in the biodiversity of boreal forests as discussed in Sec. 1. Using Optech Titan data, IBL-BalancedRF-ML achieved a recall of 64.1% and a precision of 50.6% for aspen. See B. for a full list of confusion matrices for the methods not included in Figs. 7(a)–(d).
@@ 4.2.1 Scaling laws with respect to training set size
Figure 10: Species-wise classification errors of (a) FGI-PointTransformer-DL-3D and (b) FGI-RF-ML as a function of the number of training segments in logarithmic scale for HeliALS data. The circle markers denote the number of training segments per species in the benchmark competition. For the DL method, the training set was further divided into training and validation sets for hyperparameter tuning. (c) and (d) show the normalized confusion matrices of FGI-PointTransformer-DL-3D (macro $F_{1}=0.86$) and FGI-RF-ML (macro $F_{1}=0.73$) for HeliALS data when using 5-fold cross-validation on the full ground-truth species dataset (5000 training segments).
In this section, we study the effect of the training set size on the species classification accuracy using FGI-PointTransformer-DL-3D as a representative deep learning method and FGI-RF-ML as a representative machine learning method. Here, we consider stratified subsets of the benchmarking dataset with a varying number of training tree segments ranging from 250 to 5000 as explained in Sec. 2.4. For a given subset, we utilize 5-fold cross-validation to estimate the overall and macro-average accuracy of the species classification together with the corresponding classification errors. For the scaling law studies, we excluded 88 segments from the original training set of the benchmarking competition due to non-optimum profile images, sparse point clouds, or building intersection within the segment boundaries. In addition, the species class of 18 segments from the original training set was corrected for the scaling law experiments because of errors detected during on-site verification of the dataset. However, these modifications of the dataset did not significantly affect the results of the species classification. For example, FGI-PointTransformer-DL-3D obtained an overall accuracy of 87.9% when trained on the original training set of 1065 segments in the benchmark competition, whereas the same model obtained an overall accuracy of 88.5% when trained on 1000 segments in the scaling law studies.
We illustrate the overall and macro-average accuracy of the studied models as a function of the training set size for HeliALS and Optech Titan datasets in Figs. 9(a) and (c), respectively. For the HeliALS data, we observe that the deep learning model outperforms the machine learning model already at the training set size of 250 segments. As the number of segments in the training set increases, the difference in the performance of the DL and ML models widens up. For the largest considered training set sizes, the macro-average accuracy of the deep learning method nearly catches up with the overall accuracy of the machine learning method. For a training set size of 5000 segments, FGI-PointTransformer-DL-3D reached an overall (macro-average) accuracy of 92.0% (85.1%), whereas FGI-RF-ML achieved an overall (macro-average) accuracy of 86.8% (69.6%).
@@
To gain more quantitative insight into the scaling behavior of species classification, we show the overall and macro-average classification error as a function of the training set size in a logarithmic scale in Figs. 9(b) and (d) for the HeliALS and Optech Titan data, respectively. Similarly to neural scaling laws for language and vision tasks (Hestness et al., 2017; Kaplan et al., 2020; Bahri et al., 2024), we observe a power law scaling of the classification error $\varepsilon$ following the model
@@
where $A$ is a multiplier, $m$ denotes the training set size, and $\alpha$ describes the speed of convergence. We observe that the power law scaling applies both for the point transformer model representing deep learning methods and the RF classifier representing machine learning methods. For the HeliALS data in Fig. 9(b), we notice no clear deviation from the power law fit across the studied range of training set sizes spanning an order of magnitude. For the Optech Titan dataset, the macro-average classification error of the ML model appears to slowly deviate from the power law behavior at large training set sizes, though even larger training set sizes would be needed to verify this observation. Importantly, the classification error of the DL method converges approximately twice faster with increasing training set size compared to the ML method. For example on the HeliALS dataset, the DL method is characterized by $\alpha=0.256$ ($\alpha=0.353$) regarding the overall (macro-average) error, whereas the ML method reaches $\alpha=0.14$ ($\alpha=0.159$). Moreover, we observe that the DL method has a higher $\alpha$ and converges faster on the HeliALS dataset than on the Optech Titan dataset, implying that DL methods may provide a larger advantage on dense, high-quality ALS point clouds.
@@
The power law behavior indicates that the classification accuracy may be further improved by increasing the training set size. According to our results, the benefits of deep learning methods become more apparent on large datasets consisting of thousands of trees, whereas machine learning methods may provide sufficient performance on small datasets of up to a few hundred trees. Importantly, the improvements obtained by increasing the training set size obey a law of diminishing returns. Namely, the relative reduction of the error obtained by increasing the training set size from 1000 trees to 10000 trees is expected to be equivalent to the reduction obtained by increasing the training set size from 100 to 1000 trees. Extrapolating the fitted scaling laws on the HeliALS data, we estimate that reaching a macro-average accuracy above 90% across the nine species would require 14 000 training segments using the DL method and up to $4.9\times 10^{6}$ segments using the ML approach, highlighting the advantages of deep learning on large datasets. For the Optech Titan data, we similarly estimate that $12\times 10^{6}$ segments would be required to reach a macro-average accuracy above 90% with the deep learning method. However, these estimates should be regarded as rough order of magnitude estimates, and we note that the transferability of the deep learning methods should be studied further as discussed in Sec. 4.4. The good fit of the power law model also highlights the high quality of our field reference dataset, namely, any potential classification errors in our field reference dataset must be significantly rarer than the lowest achieved error of 8.0% since we observe no clear deviation from the power law scaling.

### v2

@@ 4 Results and Discussion
In this section, we evaluate the performance of the studied species classification methods on the multispectral data collected with the HeliALS and Optech Titan systems. First in Sec. 4.1, we present the results of the benchmarking contest and discuss key factors affecting the classification accuracy, such as the use of a deep learning or a machine learning model, model architecture and training strategy. In Sec. 4.2, we analyze the effect of multispectral information on classification accuracy. Sec. 4.3 discusses the relation between classification accuracy and species. In Sec. 4.4, we analyze the scaling of classification accuracy for selected deep learning and machine learning models as a function of training set size and point density to study the impact of these factors in more detail and to highlight the differences of deep learning and machine learning methods. Subsequently, we investigate the effect of segmentation quality and crown class on species classification in Sec. 4.5. Finally, we provide further discussion and compare our results to previous studies in Sec. 4.6.
@@ 4.1 International benchmarking of species classification methods
Figure 6: (a) Comparison of overall (filled bar) and macro-average (outline) accuracy for methods using HeliALS data, with point-based DL methods shown with dark blue color, image-based DL methods shown with light blue color, and machine learning methods shown with red color. The error bars represent 95% confidence intervals based on a bootstrapping approach, see Sec. $3.$ SA-DetailView-DL-2D was not retrained on the training set of the current study, and the model parameters were directly transferred from puliti2025benchmarking. When neglecting aspens, rowans and alders not present in the FOR-species20K dataset (puliti2025benchmarking), SA-Detailview-DL-2D reached an overall accuracy of 58.8% and a macro-average accuracy of 45.4%. (b) Same as (a) but for the methods using the Optech Titan dataset.
@@
The species classification methods introduced in Sec. 3.2 were benchmarked using overall accuracy, macro-average accuracy, species-wise precision and recall as the key metrics. We compare the overall and macro-average accuracy for the methods using the HeliALS data in Fig. 6(a) and for the methods using the Optech Titan data in Fig. 6(b). In general, the methods using the HeliALS data reached a higher overall and macro-average accuracy compared to methods using the Optech Titan data, which can be attributed to higher point density and geometric accuracy of the point cloud provided by the HeliALS system, see Sec. 2.2.
@@
For the HeliALS dataset, FGI-PointTransformer-DL-3D reached the highest overall accuracy of 87.9%, and the corresponding method with a weighted loss function, FGI-PointTransformerWeighted-DL-3D, achieved the highest macro-average accuracy of 74.5%. In addition to the point transformer model, the other best performing methods on the HeliALS dataset were also point-based deep learning methods, with FGI-DGCNN-DL-3D and FGI-PointNet-DL-3D reaching an overall (macro-average) accuracy of 87.0% (67.3%) and 85.9% (64.2%), respectively. The best image-based deep learning method, DetailView-DL-2D, reached an overall (macro-average) accuracy of 84.3% (63.9%) and the only machine learning model, FGI-RF-ML, had an overall (macro-average) accuracy of 83.2% (61.3%). Based on the confidence intervals from bootstrapping, the overall accuracy of FGI-PointTransformer-DL-3D was within the confidence intervals of FGI-PointTransformerWeighted-DL-3D and FGI-DGCNN-DL-3D, suggesting that the difference between these methods is not statistically significant. The total width of the 95% confidence intervals for the overall accuracy ranged between 1.7% and 2.7% across the methods using the HeliALS dataset.
For most of the methods, the macro-average accuracy was significantly lower than the overall accuracy due to the inherent difficulty of predicting the minority species, such as alder or oak. The point transformer model clearly outperformed the other methods in the prediction of the minority species, which resulted in a significant improvement of the macro-average accuracy, especially when using the weighted loss function. FGI-PointTransformerWeighted-DL-3D achieved the highest macro-average accuracy even when taking the finite confidence intervals into account. In general, the total width of the 95% confidence intervals for the macro-average accuracy was 3.1-4.7% for the methods using the HeliALS data. Since the macro-average accuracy weighs each class equally, rare classes contribute a higher variance, leading to a wider confidence interval compared to overall accuracy.
@@
Figure 7: (a) Normalized confusion matrix of FGI-RF-ML obtaining the highest macro $F_{1}$ score (0.64) among the maching learning models using HeliALS data. FGI-RF-ML produced 55 missing predictions. (b) Normalized confusion matrix of FGI-PointTransformerWeighted-DL-3D obtaining the highest macro $F_{1}$ score (0.73) among the deep learning models using HeliALS data. (c) Normalized confusion matrix of IBL-BalancedRF-ML obtaining the highest macro $F_{1}$ score (0.57) among the machine learning models using Optech Titan data. (d) Normalized confusion matrix of FGI-PointTransformer-DL-3D obtaining the highest macro $F_{1}$ score (0.58) among the deep learning models using Optech Titan data. All confusion matrices were row-normalized to enable comparison across species regardless of the class size. Each cell count was divided by the sum of its true-class row, and thus, each row sums to 1. Diagonal values represent species-wise recalls. Precisions are displayed below the matrix. Row and column sums are shown to the right and bottom of the matrix in gray.
@@
For the Optech Titan dataset, the highest classification accuracy was reached by a random forest classifier, IBL-BalancedRF-ML, which achieved an overall accuracy of 79.9% and a macro-average accuracy of 57.6%. The point transformer method, FGI-PointTransformer-DL-3D, ranked second with an overall accuracy of 79.6% and a macro-average accuracy of 56.0%. For a larger training set size, the deep learning method is expected to outperform the machine learning model as discussed in Sec. 4.4. Based on the confidence intervals from bootstrapping, the overall accuracy of IBL-BalancedRF-ML was within the confidence intervals of FGI-PointTransformer-DL-3D, UEF-RF-ML, and FGI-PointNet-DL-3D, implying that the difference between these methods is not statistically significant. The total width of the 95% confidence intervals ranged between 2.1-2.4% for the overall accuracy and between 3.6-4.2% for the macro-average accuracy across the methods using the Optech Titan dataset.
On the Optech Titan dataset, most of the participating methods were RF classifiers or their variants. The best and worst-performing RF models had a difference of 7.9 (10.3) percentage points in terms of the overall (macro-average) accuracy implying that the performance of the various RF approaches was relatively consistent. This suggests that RF is a robust and reliable method, providing reasonable accuracy across different implementations. Based on the submissions by UEF, a RF classifier also outperformed both LGBM and SVM classifiers, though the difference was only a few percentage points.
@@
It is difficult to draw exhaustive conclusions on the factors causing the performance differences between the studied RF classifiers. However, the method descriptions in Sec. 3.2.3 suggest that the best-performing RF classifiers often utilized multi-channel intensity and echo features as well as various histogram features, such as intensity histograms. Interestingly, intensity features appear to be more important than geometric features since the second best RF method, UEF-RF-ML, only utilized a single geometric feature, but still managed to reach a good performance thanks to various intensity and echo features. When it comes to training strategies, the best performing model, IBL-BalancedRF-ML, cleaned the training dataset by removing under- or over-segmented trees, and applied a balanced RF procedure (chen2004using) to up-sample the number of training segments for minority species. However, LUKE-MultiRF-ML and Aalto-RF-ML also applied strategies to mitigate class imbalance but they were outperformed by several methods that did not use any strategy for class imbalance mitigation.
Some RF methods, such as FGI-RF-ML, had trouble evaluating all of their features on certain segments due to missing point attributes for some of the channels, thus hampering their prediction. On sparse tree segments, the computation of complicated geometric features may also not always be possible, which may lead to further missing feature values. Thus, sufficiently simple geometric features should be preferred on sparse ALS data. Imputation strategies may also be useful if the point attributes are missing for some of the channels on certain tree segments due to differences in the sensing directions of the different channels.
@@
Figure 8: Precision and recall for the best machine learning and deep learning methods for HeliALS (panels (a) and (b)) and Optech Titan (panels (c) and (d)) datasets. Best methods were chosen based on the macro-average $F_{1}$ score. Median values for precision and recall (dashed horizontal lines) were calculated across the methods utilizing the same dataset and approach (DL/ML). (a) Precision and recall for FGI-RF-ML utilizing HeliALS data (Macro $F_{1}=0.64$) (colored bars). (b) Precision and recall for the best DL method, FGI-PointTransformerWeighted-DL-3D, (Macro $F_{1}=0.73$) (colored bars) and the median across other DL methods (dashed lines) on the HeliALS dataset. (c) Precision and recall for the best ML method (Macro $F_{1}=0.57$) (colored bars) and the median across other ML methods (dashed lines) on the Optech Titan data. (d) Precision and recall for the best DL method (Macro $F_{1}=0.58$) (colored bars) and the median across other DL methods on the Optech Titan data (dashed lines).
@@
Using the HeliALS data, we also tested a sensor-agnostic method, SA-Detailview-DL-2D. The method had been previously trained on the FOR-species20K dataset  (puliti2025benchmarking), and we did not re-train it on the current dataset as explained in Sec. 3.2.2. The method did not utilize any spectral information either. On the full test set, SA-Detailview-DL-2D achieved an overall (macro-average) accuracy of 51.4% (30.3%) whereas the overall (macro-average) accuracy was 58.8% (45.4%) when neglecting aspen, rowan, and alder not present in the FOR-species20K dataset. However, the performance of the DetailView model was greatly improved by training the model from scratch using the training dataset of our study. This resulted in an overall (macro-average) accuracy of 84.3% (63.9%) for DetailView-DL-2D. The finding suggests that the model architecture itself is promising for species classification, but the model trained on the global FOR-species20K dataset did not generalize well enough for our ALS dataset. Possible reasons behind the poor out-of-domain performance of SA-Detailview-DL-2D are the relatively small proportion of high-density ALS data compared to TLS and MLS data in the FOR-species20K training dataset, the impact of segmentation artifacts, and the poor representativeness of the species of this study in the FOR-species20K training dataset.
4.2 Effect of multispectral information
@@
4.3 Species-wise analysis
To understand differences in the classification accuracy between different species, we present confusion matrices of the best-performing deep learning and machine learning models for both the HeliALS and Optech Titan datasets in Figs. 7(a)–(d). See Fig. D.1 in D for the corresponding unnormalized confusion matrices. Furthermore, Figures 8(a)–(d) illustrate the species-wise precision and recall for the best methods on both datasets together with the median of the machine learning or deep learning approaches. Based on Figs. 7(a)–(d), we observe that pines, spruces, birches, and maples are reliably classified on both datasets across the models. The other deciduous trees representing minority species (see Fig. 3) are associated with significantly lower species-wise recall and precision. Based on the confusion matrices, we observe that aspens are often classified as birches, rowans as maples, lindens as maples, alders as aspens, and oaks as all other deciduous trees across the different models and datasets. The difficulty in predicting these species arises partly due to their low frequency in the benchmarking dataset and partly due to a rather similar visual appearance as further discussed in Sec. 4.6. For the Optech Titan dataset, we observe no clear qualitative difference between the predictions of the best deep learning and the best machine learning model.
The dense HeliALS dataset enables a more accurate classification of the minority species. Importantly, the best deep learning method, FGI-PointTransformerWeighted-DL-3D, clearly outperforms the best machine learning method FGI-RF-ML in the prediction of minority species on the HeliALS dataset, which highlights the promise of deep learning for species classification on dense ALS point clouds. FGI-PointTransformerWeighted-DL-3D achieved a recall of 76.9% and a precision of 66.3% for aspen that has an important role in the biodiversity of boreal forests as discussed in Sec. 1. Using Optech Titan data, IBL-BalancedRF-ML achieved a recall of 64.1% and a precision of 50.6% for aspen. See D for a full list of normalized confusion matrices for the methods not included in Figs. 7(a)–(d).
@@ 4.4.1 Scaling laws with respect to training set size
Figure 10: Species-wise classification errors of (a) FGI-PointTransformer-DL-3D and (b) FGI-RF-ML as a function of the number of training segments in logarithmic scale for HeliALS data. The circle markers denote the number of training segments per species in the benchmark competition. For the DL method, the training set was further divided into training and validation sets for hyperparameter tuning. (c) and (d) show the row-normalized confusion matrices of FGI-PointTransformer-DL-3D (macro $F_{1}=0.86$) and FGI-RF-ML (macro $F_{1}=0.73$) for HeliALS data when using 5-fold cross-validation on the full ground-truth species dataset (5000 training segments).
In this section, we study the effect of the training set size on the species classification accuracy using FGI-PointTransformer-DL-3D as a representative deep learning method and FGI-RF-ML as a representative machine learning method. See Sec. 3.4 for details on the implementation of the scaling analysis. We illustrate the overall and macro-average accuracy of the studied models as a function of the training set size for HeliALS and Optech Titan datasets in Figs. 9(a) and (c), respectively. For the HeliALS data, we observe that the deep learning model outperforms the machine learning model already at the training set size of 250 segments. As the number of segments in the training set increases, the difference in the performance of the DL and ML models widens up. For the largest considered training set sizes, the macro-average accuracy of the deep learning method nearly catches up with the overall accuracy of the machine learning method. For a training set size of 5000 segments, FGI-PointTransformer-DL-3D reached an overall (macro-average) accuracy of 92.0% (85.1%), whereas FGI-RF-ML achieved an overall (macro-average) accuracy of 86.8% (69.6%).
@@
To gain more quantitative insight into the scaling behavior of species classification, we show the overall and macro-average classification error as a function of the training set size in a logarithmic scale in Figs. 9(b) and (d) for the HeliALS and Optech Titan data, respectively. Similarly to neural scaling laws for language and vision tasks (hestness2017deep; kaplan2020scaling; bahri2024explaining), we observe a power law scaling of the classification error $\varepsilon$ following the model
@@
where $A$ is a multiplier, $m$ denotes the training set size, and $\alpha$ describes the speed of convergence. We observe that the power law scaling applies both for the point transformer model representing deep learning methods and the RF classifier representing machine learning methods. For the HeliALS data in Fig. 9(b), we notice no clear deviation from the power law fit across the studied range of training set sizes spanning an order of magnitude. For the Optech Titan dataset, the macro-average classification error of the ML model appears to slowly deviate from the power law behavior at large training set sizes, though even larger training set sizes would be needed to verify this observation. Importantly, the classification error of the DL method converges approximately twice faster with increasing training set size compared to the ML method. For example on the HeliALS dataset, the DL method is characterized by $\alpha=0.256$ ($\alpha=0.353$) regarding the overall (macro-average) error, whereas the ML method reaches $\alpha=0.14$ ($\alpha=0.159$). Moreover, we observe that the DL method has a higher $\alpha$ and converges faster on the HeliALS dataset than on the Optech Titan dataset, implying that DL methods may provide a larger advantage on dense ALS point clouds with high geometric accuracy.
@@
The power law behavior indicates that the classification accuracy may be further improved by increasing the training set size. According to our results, the benefits of deep learning methods become more apparent on large datasets consisting of thousands of trees, whereas machine learning methods may provide sufficient performance on small datasets of up to a few hundred trees. However, the improvements obtained by increasing the training set size obey a law of diminishing returns. Namely, the relative reduction of the error obtained by increasing the training set size from 1000 segments to 10000 segments is expected to be equivalent to the reduction obtained by increasing the training set size from 100 to 1000 segments. Extrapolating the fitted scaling laws on the HeliALS data, we estimate that reaching a macro-average accuracy above 90% across the nine species would require 14 000 training segments using the DL method and up to $4.9\times 10^{6}$ segments using the ML approach, highlighting the advantages of deep learning on large datasets. For the Optech Titan data, we similarly estimate that $12\times 10^{6}$ segments would be required to reach a macro-average accuracy above 90% with the deep learning method. However, these estimates should be regarded as rough order of magnitude estimates, and we note that the transferability of the deep learning methods should be studied further as discussed in Sec. 4.6. The good fit of the power law model also highlights the high quality of our field reference dataset, namely, any potential classification errors in our field reference dataset must be significantly rarer than the lowest achieved error of 8.0% since we observe no clear deviation from the power law scaling.